Brandi Co. has a beta of 1.00. The firm currently has 30% debt, but is considering changing its capital structure to be 20% debt and 80% equity. If its corporate tax rate is 40%, what is Brandi's levered beta at 20% debt level?

Hint: First calculate unlevered beta using old capital structure and then calculate levered beta using the new capital structure.

a.

1.19

b.

1.01

c.

0.91

d.

1.10

e.

1.75
Answer #1

Let BU be unlevered beta and BL be levered beta.

First lets find unlevered beta in the present situation

BU = BL / [1 + ((1 – Tax Rate) x Debt/Equity)] = 1/(1+(1-40%)*(30%/70%))=0.7954

Now capital structure changes to 20% debt and 80% equity

BL=BU*[1 + ((1 – Tax Rate) x Debt/Equity)]=0.7954*[1 + ((1 – 40%) x (20%/80%))]=0.91 (option c)